calculate the binding energy per atoms, per nucleon, and per mole for a maganese-55 atom measured mass 54.938047 u
1
Expert's answer
2020-12-01T09:08:47-0500

Binding energy = Mass defect × c²


Mass defect = mass of the element - observed mass


Mass defect = (25×1.00728 + 30×1.00867) - 54.938047 =


0.504053 amu


Converting the mass from kg to amu,


0.504053 amu = 0.504053 × 1.66×10-²⁷ kg = 8.367 × 10-²⁸kg


E = mc² = 8.367 × 10-²⁸ × (3 × 10⁸)² = 7.53×10-¹¹J per atom


= 7.53×10-¹⁴lkj/atom



To convert to kj/mol, we multiply the energy by the number of atoms in a mole;

E = 7.53×10-¹⁴ × Avogadro's number

E = 7.53×10-¹⁴ × 6.02 × 10²³ =

4.53 × 10¹⁰kj/mole